Calculate the modulus and unit vector of
1 r1=2i-5j+7k
2 r2=3i+8j-k
1
Expert's answer
2021-07-26T15:43:19-0400

1) "|r_1|=\sqrt{2^2+(-5)^2+7^2}=\sqrt{78}."

"u_1=\frac{2}{\sqrt{78}}i-\frac{5}{\sqrt{78}}j+\frac{7}{\sqrt{78}}k"


2) "|r_2|=\sqrt{3^2+8^2+(-1)^2}=\sqrt{74}" .

"u_2=\frac{3}{\sqrt{74}}i+\frac{8}{\sqrt{74}}j-\frac{1}{\sqrt{74}}k"